# Delayed Acceptance Markov Chain Monte Carlo for Robust Bayesian Analysis

### v1

@@ Abstract
This study introduces a computationally efficient algorithm, delayed acceptance Markov chain Monte Carlo (DA–MCMC), designed to improve posterior simulation in quasi-Bayesian inference. Quasi-Bayesian methods, which do not require fully specifying a probabilistic model, are often computationally expensive owing to the need to evaluate the inverse and determinant of large covariance matrices. DA–MCMC addresses this challenge by employing a two-stage process: In the first stage, proposals are screened using an approximate posterior, whereas a final acceptance or rejection decision is made in the second stage based on the exact target posterior. This reduces the need for costly matrix computations, thereby improving efficiency without sacrificing accuracy. We demonstrate the effectiveness of DA–MCMC through applications to both synthetic and real data. The results demonstrate that, although DA–MCMC slightly reduces the effective sample size per iteration compared with the standard MCMC, it achieves substantial improvement in terms of effective sample size per second, approximately doubling the efficiency. This makes DA–MCMC particularly useful for cases where posterior simulation is computationally intensive. Thus, the DA–MCMC algorithm offers a significant advancement in computational efficiency for quasi-Bayesian inference, making it a valuable tool for robust Bayesian analysis.
Keywords: delayed acceptance Markov chain Monte Carlo; robust Bayesian analysis; generalized posterior; Gibbs posterior
@@ 2 Delayed acceptance MCMC
To expedite the posterior simulation with (2), we propose using DA–MCMC [6]. A single cycle of the algorithm comprises two stages: (1) screening a proposal with an approximate posterior and (2) determining whether to accept or reject with the exact posterior. The second stage occurs only when a proposal passes the first stage.
Given a current state $\boldsymbol{\theta}^{\left(t\right)}$, a proposal $\boldsymbol{\theta}^{\prime}$ is generated from a proposal kernel, $\boldsymbol{\theta}^{\prime}\sim q_{1}\left(\cdot|\boldsymbol{\theta}^{\left(t\right)}\right)$. The first-stage acceptance probability is based on an approximation to the exact target kernel $\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)$, which is evaluated with less computational cost than $\pi\left(\boldsymbol{\theta}\right)$. An approximation to the target kernel is constructed by replacing $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)$ with the current value $\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)$:
$\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)=\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)\right|^{-\frac{1}{2}}\exp\left\{-\frac{N}{2}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}^{\prime}\right)^{\top}\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)^{-1}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}^{\prime}\right)\right\}p\left(\boldsymbol{\theta}^{\prime}\right).$ (4)
With this approximation, we do not need to compute the most computationally demanding quantities, that is, the inverse and determinant of $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)$. The first-stage acceptance probability is computed as follows:
$\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{1}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)}{q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi^{*}\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\}.$ (5)
With probability $\alpha_{1}\left(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta}^{\left(t\right)}\right)$, $\boldsymbol{\theta}^{\prime}$ is “promoted” as a proposal for the second stage; otherwise, $\boldsymbol{\theta}^{\prime}$ is rejected, thereby maintaining the current state $\boldsymbol{\theta}^{\left(t+1\right)}=\boldsymbol{\theta}^{\left(t\right)}$. In the second stage, the acceptance probability is evaluated based on the exact target kernel $\pi\left(\boldsymbol{\theta}^{\prime}\right)$ as follows:
$\alpha_{2}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{2}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi\left(\boldsymbol{\theta}^{\prime}\right)}{q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\}.$ (6)
where $q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)$ denotes the actual proposal density given by:
$q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right).$ (7)
@@
As argued in [6], compared with the standard MCMC, a single cycle of DA–MCMC can be slightly less efficient in terms of effective sample size per iteration. However, DA–MCMC can be more efficient in terms of the effective sample size per second because it does not have to perform the unnecessary computation of $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)^{-1}$ and $\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)\right|$. As shown in the next section, DA–MCMC is effective in our context.
@@
This study uses an adaptive random walk Metropolis–Hastings kernel, $q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\mathrm{N}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)},\;\varepsilon\boldsymbol{\Sigma}\right)$, where $\varepsilon>0$ denotes a step size, $\boldsymbol{\Sigma}$ denotes the $K$-dimensional symmetric positive definite matrix, and $\mathrm{N}\left(\boldsymbol{x}|\boldsymbol{a},\boldsymbol{B}\right)$ denotes a probability density function of a multivariate normal distribution with mean $\boldsymbol{a}$ and covariance matrix $\boldsymbol{B}$ evaluated at $\boldsymbol{x}$. The parameters of the proposal kernel are adaptively selected on the fly of an MCMC algorithm [8]. At the $t$-th cycle, $\varepsilon$ is updated as $\log\varepsilon\leftarrow\log\varepsilon+t^{-\varsigma}\left(\bar{\alpha}^{\left(t\right)}-\alpha^{\dagger}\right)$, where $\bar{\alpha}^{\left(t\right)}$ denotes the average acceptance probability up to the $t$-th cycle, $\alpha^{\dagger}\in\left(0,1\right)$ denotes the target acceptance probability, and $\varsigma\in\left(0.5,1\right)$ denotes a tuning parameter. In this study, we fix $\varsigma=0.51$. $\boldsymbol{\Sigma}$ is updated based on the empirical covariance matrix of $\boldsymbol{\theta}^{\left(1\right)},\boldsymbol{\theta}^{\left(2\right)},...,\boldsymbol{\theta}^{\left(t\right)}$. The adaptation of $\varepsilon$ and $\boldsymbol{\Sigma}$ is performed only in the warmup phase and not thereafter.
@@ Acknowledgements
This work was supported by JSPS KAKENHI Grant Number 20K22096.

### v2

@@ Abstract
This study introduces a computationally efficient algorithm, delayed acceptance Markov chain Monte Carlo (DA–MCMC), designed to improve posterior simulation in quasi-Bayesian inference. Quasi-Bayesian methods, which do not require fully specifying a probabilistic model, are often computationally expensive owing to the need to evaluate the inverse and determinant of large covariance matrices. DA–MCMC addresses this challenge by employing a two-stage process: In the first stage, proposals are screened using an approximate posterior, whereas a final acceptance or rejection decision is made in the second stage based on the exact target posterior. This reduces the need for costly matrix computations, thereby improving efficiency without sacrificing accuracy. We demonstrate the effectiveness of DA–MCMC through applications to both synthetic and real data. The results demonstrate that, although DA–MCMC slightly reduces the effective sample size per iteration compared with the standard MCMC, it achieves substantial improvement in terms of effective sample size per second, approximately doubling the efficiency. This makes DA–MCMC particularly useful for cases where posterior simulation is computationally intensive. Thus, the DA–MCMC algorithm offers a significant advancement in computational efficiency for quasi-Bayesian inference, making it a valuable tool for robust Bayesian analysis. …
keywords:
delayed acceptance Markov chain Monte Carlo; robust Bayesian analysis; generalized posterior; Gibbs posterior
@@ 2 Delayed acceptance MCMC
To reduce computational costs, this study proposes to employ a DA–MCMC algorithm [5] which proceeds in two stages. We first approximate the target posterior distribution $\pi\left(\boldsymbol{\theta}\right)$ with a computationally cheaper surrogate $\pi^{*}\left(\boldsymbol{\theta}\right)$, which avoids evaluating the most expensive components—specifically, the inverse and determinant of $\boldsymbol{W}\left(\boldsymbol{\theta}\right)$. The approximate posterior is constructed by replacing $\boldsymbol{W}\left(\boldsymbol{\theta}\right)$ with the current value $\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)$, as follows:
$\pi^{*}\left(\boldsymbol{\theta}\right)=\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)\right|^{-\frac{1}{2}}\exp\left\{-\frac{N}{2}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)^{\top}\boldsymbol{W}\left(\boldsymbol{\theta}^{\left(t\right)}\right)^{-1}\bar{\boldsymbol{m}}\left(\boldsymbol{\theta}\right)\right\}p\left(\boldsymbol{\theta}\right).$
In the first stage, a candidate $\boldsymbol{\theta}^{\prime}$ is proposed from a proposal distribution $q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)$. It is then preliminarily accepted with probability:
$\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{1}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi^{*}\left(\boldsymbol{\theta}^{\prime}\right)}{q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi^{*}\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\}.$
If accepted, the candidate is promoted to the second stage, where it is evaluated under the full posterior $\pi\left(\boldsymbol{\theta}\right)$. Otherwise, it is rejected and the chain remains at the current state $\boldsymbol{\theta}^{\left(t+1\right)}=\boldsymbol{\theta}^{\left(t\right)}$.
In the second stage, the final acceptance probability is computed as:
$\alpha_{2}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)=\min\left\{1,\;\frac{q_{2}\left(\boldsymbol{\theta}^{\left(t\right)}|\boldsymbol{\theta}^{\prime}\right)\pi\left(\boldsymbol{\theta}^{\prime}\right)}{q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)\pi\left(\boldsymbol{\theta}^{\left(t\right)}\right)}\right\},$
where the effective second-stage proposal distribution is corrected as:
$q_{2}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\alpha_{1}\left(\boldsymbol{\theta}^{\left(t\right)},\boldsymbol{\theta}^{\prime}\right)q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right).$
This structure improves computational efficiency by rejecting implausible proposals early, while still ensuring exact sampling from the full posterior distribution.
@@
As argued in [5], compared with the standard MCMC, a single cycle of DA–MCMC is slightly less efficient in terms of effective sample size per iteration. However, DA–MCMC can be more efficient in terms of the effective sample size per second because it does not require the unnecessary computation of $\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)^{-1}$ and $\left|\boldsymbol{W}\left(\boldsymbol{\theta}^{\prime}\right)\right|$. As shown in the next section, DA–MCMC is effective in our context.
This study uses an adaptive random walk Metropolis–Hastings kernel specified as:
$q_{1}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)}\right)=\mathrm{N}\left(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{\left(t\right)},\;\varepsilon\boldsymbol{\Sigma}\right),$
@@
where $\varepsilon>0$ denotes the step size, $\boldsymbol{\Sigma}$ denotes the $K$-dimensional symmetric positive definite matrix, and $\mathrm{N}\left(\boldsymbol{x}|\boldsymbol{a},\boldsymbol{B}\right)$ denotes the probability density function of a multivariate normal distribution with mean $\boldsymbol{a}$ and covariance matrix $\boldsymbol{B}$ evaluated at $\boldsymbol{x}$. The parameters of the proposal kernel are adaptively selected on the fly of the MCMC algorithm [7]. At the $t$-th cycle, $\varepsilon$ is updated as $\log\varepsilon\leftarrow\log\varepsilon+t^{-\varsigma}\left(\bar{\alpha}^{\left(t\right)}-\alpha^{\dagger}\right)$, where $\bar{\alpha}^{\left(t\right)}$ denotes the average acceptance probability up to the $t$-th cycle, $\alpha^{\dagger}\in\left(0,1\right)$ denotes the target acceptance probability, and $\varsigma\in\left(0.5,1\right)$ denotes the tuning parameter. In this study, we fix $\varsigma=0.51$. $\boldsymbol{\Sigma}$ is updated based on the empirical covariance matrix of $\boldsymbol{\theta}^{\left(1\right)},\boldsymbol{\theta}^{\left(2\right)},...,\boldsymbol{\theta}^{\left(t\right)}$. The adaptation of $\varepsilon$ and $\boldsymbol{\Sigma}$ is performed only during the warmup phase and not thereafter.
@@ Acknowledgements
This work was supported by JSPS KAKENHI Grant Numbers 20K22096 and 25K21168.